OpenStudy (anonymous):

can someone rationalize this? sqrt2 - 1 all over sqrt 6

OpenStudy (anonymous):

Okay I would advise you to Multiply and Divide by sqrt2 + 1

OpenStudy (anonymous):

\[(\sqrt{2}-1/\sqrt{6}) * (\sqrt{6} /\sqrt{6}) \]

OpenStudy (anonymous):

Why ishaan? there is only sqrt6 in the denominator

OpenStudy (anonymous):

ah understood the question wrong sorry

OpenStudy (anonymous):

no worries

OpenStudy (anonymous):

I took denominator as sqrt2 - 1

OpenStudy (anonymous):

so do i multipy top and bottom by sqrt 6?

OpenStudy (anonymous):

Lammy, can you multiply that? give it a shot

OpenStudy (anonymous):

yes, that will eliminate the sqrt in the denominator

OpenStudy (anonymous):

ok got it so rationalize is just geting rid of sqrt on the denominator

OpenStudy (anonymous):

yes also, your instructor will probably want you to simplify that numerator after you distribute the sqrt6 to the numerator.. so keep that in mind as well

OpenStudy (anonymous):

yea = 2sqrt3-sqrt6 all over 6

OpenStudy (anonymous):

perfect, well done

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

anytime

OpenStudy (amistre64):

ishaan was right, but the question was wrong ;)